OpenStudy (anonymous):

factor -2k - k^3 - 3k^2

OpenStudy (anonymous):

-k(k^2 + 3k + 2) = -k(k +1)(k + 2)

OpenStudy (anonymous):

awesome, thanks... can you help me with this one? The volume of a rectangular prism is equal to x3 + 4x2 + 3x. What is the longest edge of the prism?

OpenStudy (anonymous):

*The volume of a rectangular prism is equal to x^3 + 4x^2 + 3x. What is the longest edge of the prism?*

OpenStudy (anonymous):

Since x^3 +4x^2 +3x = x( + 1)(x + 3), and if the dimensions are integers, then x + 3 is the longest edge.